Consider an object moving along a line with the following velocity and initial position. Assume time t is measured in seconds and velocities have units of​ m/s. Complete parts​ (a) through​ (d) below.Consider an object moving along a line with the following velocity and initial position. Assume time t is measured in seconds
Answer #1

Page - 1 vlt) = -1 -2 Ars (216 asfert) 1 1 V + -re v is so ottel ve the for 1 st 23/ (Ezo] (t = 1 = 1/2 Vlt) = ds 1 - 2 her)